OpenStudy (anonymous):

Multiply: (x^(2) -4)/x^(2)) * (x^(2)-2x)/x^(2)+3x-10

OpenStudy (anonymous):

\[\frac{ x ^{2} -4}{ x ^{2} } * \frac{ x^{2} - 2x }{ x^{2}+3x-10 }\]

OpenStudy (anonymous):

So I factored it out to be \[\frac{ (x-2)(x+2) }{ x(x) }*\frac{ x(x-2) }{ (x+5)(x-2)}\]

OpenStudy (anonymous):

And then I divided out to be \[\frac{ x+2 }{ x(x+5)}\] and that x cannot equal -5 and 2, but my homework is saying that's not correct. Can someone help me out? @phi @dan815

OpenStudy (phi):

you have (x-2) in the top twice. You can cancel one of them with the bottom (x-2), but that still leaves one up top

OpenStudy (anonymous):

Alright, that's what I thought. is that all that I did wrong?

OpenStudy (phi):

yes, all you need is multiply by (x-2) (up top)

OpenStudy (anonymous):

So, my final product would look like {x+2 (x-2)/x(x+5)}

OpenStudy (anonymous):

With x not being able to equal -5 or 2, and those two numbers only?

OpenStudy (anonymous):

Okay so I got the question wrong because I did not include 0 as a number x cannot equal. Why would I need to include that if it is obvious and I did not need to include it on my other problems?

OpenStudy (phi):

maybe because you canceled out an x. Technically you have a "double 0"

OpenStudy (anonymous):

Can you explain that concept to me? I don't remember hearing it.

OpenStudy (phi):

I think x/x^2 with x=0 is different from 1/x with x=0 but I am not a mathematician so I have no good explanation. Just a guess as to why they want x=0 listed in the excluded values.